# Dynamical Schwinger Effect in Lead-Halide Perovskites

### v1

@@ III Discussion
As an application of this sensitivity of PL to $E_{1}$ we will use it to study the highly debated frozen-in electric fields $E_{DC}$ in LHPs Ambrosio et al. (2022), which lead to a broken inversion symmetry even for nominally cubic lead-halide perovskites. It is worth noting here that the method merits certain advantages as compared to some of the conventional techniques for detecting local breaking of inversion symmetry such as second-harmonic generation or angle-resolved photoemission. First of all, our method is not confined to the vicinity of the sample sufrace since the we only use sub-gap frequencies. Second, the peculiarity of tunneling is that it is sensitive to the absolute value of field amplitudes, cf. Eq. (1). Therefore there is a finite net effect due to local inversion breaking even if the local fields spatially average to zero. Finally, another implication of the field cooperation is that PL should strongly depend on the polarization of the external field.
Figure 2: A) PL as a function of the applied electric field for two orthogonal polarization orientations aligned with the crystal axes of MAPbBr3 (markers). The corresponding fits according to our model to the data are presented as solid curves. The inset shows a cartoon of MAPbBr3 with hypothetical ferroelectric domains overlaid with a schematic representation of the irradiating laser spot whose diameter is $d\approx 200\mu$m ($1/e^{2}$); B) Polarization scan taken at a fixed value of the external electric field $eaE_{AC}\simeq 0.180$.
For experimental validation, we measure PL from a single-crystal MAPbBr3 as a function of intensity of the driving mid-infrared $\lambda=4\mu$m radiation. The resulting dependence plotted in Fig. 2A indicates that the tunneling ionization efficiency depends strongly on the polarization of the pumping field. To elucidate the nature of this polarization dependence we scan PL as a function of polarization at a fixed pump intensity as shown in Fig. 2B.
The two-fold symmetric polarization dependence in Fig. 2B might appear surprising since this measurement was performed at room temperature where MAPbBr3 is expected to be in the cubic phase, i.e., four-fold symmetric (e.g., consider the two-photon absorption experiments like in Ref. Li et al. (2020)). The data in Fig. 2B therefore indicates an unexpected lowering of symmetry in MAPbBr3 in the nominally cubic phase. In the most straightforward explanation this could be attributed to an extrinsic factor such as residual stress in the crystal structure that would lift the four-fold symmetry of the electronic band structure. However it can be demonstrated that, that in this case in order to reproduce the magnitude of the effect in Fig. 2 the relative anisotropy of microscopic Hamiltonian describing the system has to reach unrealistic levels of $\sim 50\%$ (see Supplementary Material).
@@
In an alternative scenario, the observed anisotropy can be attributed to the local frozen-in electric fields, that were conjectured to be present in lead-halide perovskites Niesner et al. (2016); Rakita et al. (2017). Indeed the exponential sensitivity of PL on applied fields implies that one can obtain the anisotropy observed in Fig. 2 already with moderate static fields. In Figs. 2A and B we show that all of the experimental data can be reproduced within the two-field Landau-Dykhne formalism (see Supplementary Material). Based on the fitting, we estimate the numeric value for the internal static fields in room-temperature single-crystal MAPbBr3 to be of the order of $E_{DC}\sim 0.1$V/nm. Within this interpretation of the PL anisotropy, the PL gets maximum enhancement when the polarization of the driving optical field coincides with the direction of $E_{DC}$. That this direction happens to be aligned with direction of Pb-Br bonds ([001] of the cubic lattice) is consistent with previous observations of possible ferroelectricity in MAPbI3 Röhm et al. (2019); Leonhard et al. (2019).
Importantly, we also observe that the two-fold symmetric pattern of PL is subject to the spatial extent of the focal spot of the driving laser. Namely, we observe that that the two-fold symmetry in Fig. 2 is only present for sufficiently focused beams. For larger illumination spots, the pattern turns out to be four-fold symmetric, which can be naturally explained if more than one ferroelectric domain with different field directions participate in the fluorescence process (see Supplementary Material). This interpretation agrees well with the previous estimations of the size of possible ferroelectric domains of several microns in MAPbI3 Rothmann et al. (2017); Rakita et al. (2017).
@@ III.2 Cooperation of dynamic fields and AC-biasing
The hallmark feature of quantum tunneling is its exponential sensitivity to external parameters, which has allowed applications defying the classical common sense. The most celebrated example is the scanning tunneling microprobes where exponential sensitivity of the tunneling current means that the overwhelming majority of it is flowing through the single atom that happens to stick out the most toward the sample  Binnig et al. (1982); Chen (2007). Realization of this has led to unprecedented sub-atom-scale spatial resolutions. Other celebrated applications of tunneling-related process include the tunnel diode Turner (1976) with its unique I-V characteristics and high-harmonic generation in atomic optics where deeply sub-threshold laser light promotes electrons in an atom to unoccupied states Yue and Gaarde (2022), thus initializing the process of extreme frequency conversion of radiation.
@@
As a demonstration of the practical use of this effect, we present in Fig. 3E the effect of AC-biasing on PL produced by a mid-infrared beam ($4.5\mu$m) in MAPbBr3333 biased with $1\mu$m radiation at $I_{\mathrm{bias}}=5\times 10^{13}$/m2. Here, we show the differential photoluminescence for two beams $\Delta PL(I)=PL(I_{1},I_{\mathrm{bias}})-PL(I_{1},0)-PL(0,I_{\mathrm{bias}})$, where $PL(I_{1},I_{\mathrm{bias}})$ stands for the total PL emitted by the sample when irradiated by mid- and near-infrared pulses overlapping in time. As can be seen the differential sensitivity of PL can be made linear in the intensity of the mid-IR which opens broad possibilities for efficient mid-infrared sensors based upon lead-halide perovskites. We demonstrate the latter by overlapping the two beams (mid- and near-infrared) on a screen covered by LHP microparticles. As can be seen in Fig. 3F-G, there is clear enhancement in local PL from the region of spatial overlap. Curiously, due to the nonlinear sensitivity of PL on the pumping beam intensity, the apparent size of the overlap spot is significantly smaller than the spot sizes measured for both beams, which can be employed for super-resolution microscopy.
@@ IV Summary
To summarize, we have demonstrated dynamical Schwinger effect in a Dirac semiconductor (lead-halide perovskite) and quantitatively described it with the quasi-adiabatic Landau-Dykhne approach. Further, we have employed the characteristic exponential sensitivity of tunneling ionization on driving fields to measure the local frozen-in fields in the nominally cubic MAPbBr3 single crystal at room temperature contributing to the ongoing debate on the state of ferroelectricity in lead-halide perovskites. Finally, we have investigated the cooperation between two time-dependent fields simultaneously driving tunneling ionization and demonstrated this cooperation can act as an AC analogue of biasing for an optical frequency upconversion, which paves the way for a mid-infrared light detection with lead-halide perovskites.

### v2

@@ III Discussion
As an application of this sensitivity of PL to $E_{1}$ we will use it to study the highly debated frozen-in electric fields $E_{DC}$ in LHPs Ambrosio et al. (2022), which lead to a broken inversion symmetry even for nominally cubic lead-halide perovskites. It is worth noting here that the method merits certain advantages as compared to some of the conventional techniques for detecting local breaking of inversion symmetry such as second-harmonic generation or angle-resolved photoemission. First of all, our method is not confined to the vicinity of the sample surface since the used electric fields have sub-gap frequencies and hence can penetrate the sample. Second, the peculiarity of tunneling is that it is sensitive to the absolute value of field amplitudes, cf. Eq. (1). Therefore there is a finite net effect due to local inversion breaking even if the local fields spatially average to zero. Finally, another implication of the field cooperation is that PL should strongly depend on the polarization of the external field.
Figure 2: A) PL as a function of the applied electric field for two orthogonal polarization orientations aligned with the crystal axes of MAPbBr3 (markers). The corresponding fits according to our model to the data are presented as solid curves. The inset shows a cartoon of MAPbBr3 with hypothetical ferroelectric domains overlaid with a schematic representation of the irradiating laser spot whose diameter is $d\approx 200\mu$m ($1/e^{2}$); B) Polarization scan taken at a fixed value of the external electric field $eaE_{AC}\simeq 0.180$eV.
For experimental validation, we measure PL from a single-crystal MAPbBr3 as a function of intensity of the driving mid-infrared $\lambda=4\mu$m radiation. The resulting dependence plotted in Fig. 2A indicates that the tunneling ionization efficiency depends strongly on the polarization of the pumping field. To elucidate the nature of this polarization dependence, we scan PL as a function of polarization at a fixed pump intensity as shown in Fig. 2B.
The two-fold symmetric polarization dependence in Fig. 2B might appear surprising since this measurement was performed at room temperature where MAPbBr3 is expected to be in the cubic phase, i.e., four-fold symmetric (e.g., consider the two-photon absorption experiments like in Ref. Li et al. (2020)). The data in Fig. 2B therefore indicates an unexpected lowering of symmetry in MAPbBr3 in the nominally cubic phase. In the most straightforward explanation this could be attributed to an extrinsic factor such as residual stress in the crystal structure that would lift the four-fold symmetry of the electronic band structure. However it can be demonstrated that, that in this case in order to reproduce the magnitude of the effect in Fig. 2 the intrinsic stains in the system have to reach relatively large ($\simeq 0.5$%) levels Chen et al. (2020), which are not expected in a single-crystal LHP Li et al. (2024) (see Supplementary Material for more details).
@@
In an alternative scenario, the observed anisotropy can be attributed to the local frozen-in electric fields, that were conjectured to be present in lead-halide perovskites Niesner et al. (2016); Rakita et al. (2017). Indeed the exponential sensitivity of PL on applied fields implies that one can obtain the anisotropy observed in Fig. 2 already with moderate static fields. In Figs. 2A and B we show that all of the experimental data can be reproduced within the two-field Landau-Dykhne formalism presented in the Supplementary Material. Based on the fitting, we estimate the numeric value for the internal static fields in room-temperature single-crystal MAPbBr3 to be of the order of $E_{DC}\sim 0.1$V/nm. Within this interpretation of the PL anisotropy, the PL gets maximum enhancement when the polarization of the driving optical field coincides with the direction of $E_{DC}$. That this direction happens to be aligned with direction of Pb-Br bonds ([001] of the cubic lattice) is consistent with previous observations of possible ferroelectricity in MAPbI3 Röhm et al. (2019); Leonhard et al. (2019). Note that Fig. 1D is produced using electric fields that are stronger than those presented in Fig. 2A. Further, we rotated the polarization of the external electric field to have minimal PL intensity. This allowed us to minimize the effect of the frozen-in electric fields in the data demonstrated in Fig. 1D.
Importantly, we also observe that the two-fold symmetric pattern of PL is subject to the spatial extent of the focal spot of the driving laser. Namely, we observe that that the two-fold symmetry in Fig. 2 is only present for sufficiently focused beams. For larger illumination spots, the pattern turns out to be four-fold symmetric (see Supplementary Material). This can be naturally explained if more than one ferroelectric domain with different field directions participate in the fluorescence process. This interpretation agrees well with the previous estimations of the size of possible ferroelectric domains of several microns in MAPbI3 Rothmann et al. (2017); Rakita et al. (2017).
@@ III.2 Cooperation of dynamic fields and AC-biasing
The hallmark feature of quantum tunneling is its exponential sensitivity to external parameters, which has enabled applications defying the classical common sense. The most celebrated example is the scanning tunneling microprobes where exponential sensitivity of the tunneling current means that the overwhelming majority of it is flowing through the single atom that happens to stick out the most toward the sample  Binnig et al. (1982); Chen (2007). Realization of this has led to unprecedented sub-atom-scale spatial resolutions. Other celebrated applications of tunneling-related process include the tunnel diode Turner (1976) with its unique I-V characteristics and high-harmonic generation in atomic optics where deeply sub-threshold laser light promotes electrons in an atom to unoccupied states Yue and Gaarde (2022), thus initializing the process of extreme frequency conversion of radiation.
@@
As a demonstration of the practical use of this effect, we present in Fig. 3E the effect of AC-biasing on PL produced by a mid-infrared beam ($4.5\mu$m) in MAPbBr3333 biased with $1\mu$m radiation at $I_{\mathrm{bias}}=5\times 10^{13}$/m2. Here, we show the differential photoluminescence for two beams $\Delta PL(I)=PL(I_{1},I_{\mathrm{bias}})-PL(I_{1},0)-PL(0,I_{\mathrm{bias}})$, where $PL(I_{1},I_{\mathrm{bias}})$ stands for the total PL emitted by the sample when irradiated by mid- and near-infrared pulses overlapping in time. As can be seen the differential sensitivity of PL can be made linear in the intensity of the mid-IR which opens broad possibilities for efficient mid-infrared sensors based upon lead-halide perovskites. We demonstrate the latter by overlapping the two beams (mid- and near-infrared) on a screen covered by LHP microparticles. As can be seen in Fig. 3F-G, there is clear enhancement in local PL from the region of spatial overlap. Curiously, due to the nonlinear sensitivity of PL on the pumping beam intensity, the apparent size of the overlap spot is significantly smaller (11 $\mu$m) than the actual spot size of the mid-infrared pulse (160 $\mu$m), which can be employed for super-resolution microscopy.
@@ IV Summary
To summarize, we have demonstrated dynamical Schwinger effect in a Dirac semiconductor (lead-halide perovskite) and quantitatively described it with the quasi-adiabatic Landau-Dykhne approach. Further, we have employed the characteristic exponential sensitivity of tunneling ionization on driving fields to measure the local frozen-in fields in the nominally cubic MAPbBr3 single crystal at room temperature contributing to the ongoing debate on the state of ferroelectricity in lead-halide perovskites. Finally, we have investigated the cooperation between two time-dependent fields simultaneously driving tunneling ionization and demonstrated that this cooperation can act as an AC analogue of biasing for an optical frequency upconversion. These findings pave the way for a mid-infrared light detection with lead-halide perovskites.
Acknowledgements A. G. V. thanks Peter Balling for useful discussions.
Supplementary Material for ‘Dynamical Schwinger effect and non-perturbative light detection in lead halide perovskites’
Ionization in the presence of a frozen-in electric field.
Let us assume that the vector potential has the form
$\mathbf{A}=-\mathbf{E}_{DC}(t-t_{i})-\mathbf{E}_{AC}\frac{\sin(\omega t)}{\omega},$ (11)
where $\mathbf{E}_{DC}$ and $\mathbf{E}_{AC}$ are constant vectors that define the strength of the DC and AC electric fields, respectively; $t_{i}$ is some constant that is determined by initial conditions. We assume that $\mathbf{A}(t=0)=0$, which means that the difference between energy levels at $t=0$ is $\Delta$. This implies that $t_{i}=0$.
It is clear that to satisfy $\mathcal{E}_{f}(t)=\mathcal{E}_{i}(t)$, $t$ should be imaginary, i.e., $t=i\tau$, where $\tau$ satisfies the equation
$E_{DC}^{2}\tau^{2}+2(\mathbf{E}_{DC}\mathbf{E}_{AC})\frac{\tau\sinh(\omega\tau)}{\omega}+E_{AC}^{2}\frac{\sinh^{2}(\omega\tau)}{\omega^{2}}=\frac{\Delta^{2}}{(ea)^{2}(\Delta t_{3}+16t^{2})}.$ (12)
We assume that $E_{DC}\to 0$ so that
$eaE_{DC}\tau\cos(\theta)+eaE_{AC}\frac{\sinh(\omega\tau)}{\omega}\simeq\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}},$ (13)
where $\theta$ is the angle between $\mathbf{E}_{AC}$ and $\mathbf{E}_{DC}$. We can solve this equation iteratively. At the first iteration, we neglect the static field completely:
$eaE_{AC}\frac{\sinh(\omega\tau^{0})}{\omega}\simeq\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}}\to\tau^{0}=\frac{1}{\omega}\mathrm{arsinh}\left(\frac{\omega}{eaE_{AC}}\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}}\right).$ (14)
This expression reproduces the result for one-color tunnel ionization. At the second iteration, we derive:
$eaE_{DC}\tau^{0}\cos(\theta)+eaE_{AC}\frac{\sinh(\omega\tau^{1})}{\omega}\simeq\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}}\to$ (15)
$\tau^{1}=\frac{1}{\omega}\mathrm{arsinh}\left(\frac{\omega}{eaE_{AC}}\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}}-\frac{E_{DC}}{E_{AC}}\omega\tau^{0}\cos(\theta)\right).$
We have checked numerically that for $E_{DC}\ll E_{AC}$, $\tau^{1}$ approximates $\tau$ well.
The expression for $\tau_{1}$ allows us to compute $W_{fi}$:
$W_{fi}\sim\mathrm{exp}\left(-2\mathrm{Im}\int_{0}^{i\tau^{1}}\left[\Delta+\frac{\Delta t_{3}+16t^{2}}{2\Delta}\left(ea\mathbf{A}\right)^{2}\right]\mathrm{dt}\right),$ (16)
so that
$W_{fi}\sim\mathrm{exp}\left(-2\tau^{1}\Delta+(eaE_{AC})^{2}\frac{\Delta t_{3}+16t^{2}}{4\Delta}\left[\frac{\sinh(2\omega\tau_{1})-2\omega\tau_{1}}{\omega^{3}}+\frac{8E_{DC}}{E_{AC}}\cos(\theta)\frac{\omega\tau_{0}\cosh(\omega\tau_{0})-\sinh(\omega\tau_{0})}{\omega^{3}}\right]\right),$ (17)
where, for consistency of derivations, we use $\tau_{0}$ when we multiply by $E_{DC}$. Note that because the direction of $\mathbf{E}_{AC}$ is defined only up to a phase $\pi$, we should use absolute value of $\cos(\theta)$ in the expression.
Additional information for the discussion on frozen-in electric fields
As is mentioned in the main text, the two-fold symmetric polarization dependence in Fig. 2B of the main text appears surprising since the measurement is performed at room temperature where MAPbBr3 is expected to be in the cubic phase. To further understand this, we approximate the rate of tunneling ionization with the following expression
$W_{fi}\sim e^{-\frac{5}{3}\frac{\Delta}{eaE}\sqrt{\frac{\Delta^{2}}{\Delta t_{3}+16t^{2}}}},$ (18)
which is valid in an adiabatic regime. We see that there are a few possibilities to change the rate as a function of the angle. For example, either $E$ or $t_{3}$ and $t$ can depend on the angle. Let us first assume that $E$ is independent of the angle and that $t\to t+\delta t$ and $t_{3}\to t_{3}+\delta t_{3}$, where $\delta t$ and $\delta t_{3}$ determine the change of the hopping integrals. In this case, for the parameters of our system, we derive
$W_{fi}\sim e^{-1.4\frac{\Delta}{eaE}}e^{0.5\frac{\delta}{eaE}},$ (19)
where $\delta=\delta t_{3}+32t\frac{\delta t}{\Delta}\simeq\delta t_{3}+8\delta t$. Assuming that $\delta t_{3}\sim\delta t$, we have $\delta\simeq 10\delta t$. Our data show that for $eaE=0.140$eV the change in PL intensity can be as large as 40, see Fig. 2B of the main text. This means that the change of the hopping coefficient can be $\delta t\simeq 0.05$eV. Such a change implies that the effective mass is modified by about $10-20$%, which can be achieved if the intrinsic strains in a single-crystal LHPs are of the order of 0.5% Chen et al. (2020). Depending on the preparation of the LHP single crystal, it can either be strain free or have strains $\simeq 0.1$Li et al. (2024). We are not aware of a mechanism that can lead to a large renormalization of the effective mass under such weak strains. Therefore, we conclude that the electric field $E$ depends on angle, which leads to the hypothesis of ferroelectric domains presented in the main text.
[…]
Figure 4: Polarization scan taken at $eaE_{AC}\simeq 0.284$eV (red markers) and $eaE_{AC}\simeq 0.312$eV (blue markers). The PL intensity is shown in arbitrary units.
Hysteresis of PL intensity
[…]
Abraded samples of MAPbBr3
As was mentioned above, the efficiency of tunneling ionization is very sensitive to the local properties of the sample (stress, static fields, etc). Indeed, it has been observed that photoluminescence becomes visibly enhanced when the irradiated region happens to be situated near structural defects such as surface inhomogeneities or sample edges/cracks. This is unfortunate for possible infrared imaging applications of dynamical SE in perovskites, as structure-induced variations of PL are going to obscure the intrinsic PL image reflecting the spatial distribution of the irradiating mid-infrared wave-front.
Sample imperfections are unavoidable in practice. Therefore, not being able to ensure sample homogeneity by removing the imperfections one can take the other extreme and instead homogenize the defect distribution, thus obtaining a quasi-uniform sample suitable for infrared-imaging purposes.
Specifically, in this work we mechanically grind the sample by rubbing an intact single crystal of MAPbBr3 against a fine 1 µm-grit Al2O3. As a result, we obtain a sample consisting of a substrate (sandpaper) covered by a powder of perovskite microcrystals as shown in Fig. 6. Our experience shows that microcrystal samples obtained this way are very robust, retaining the ability to PL under infrared-irradiation for years after being manufactured even if stored in atmosphere. The additional advantage of such samples as compared to single-crystal ones is the former being more sensitive than the latter since local stresses as a rule tend to enhance PL efficiency.
To characterize the morphology of the microcrystal comprising the abraded samples, we first cut the sandpaper with microcrystals on top of them into small mm-sized pieces, which were then dropped into hexane and sonicated for 5 minutes. The resulting solution was drop casted on silicon substrates and allowed to dry. In this way prepared microcrystal samples were subsequently imaged by a SEM microscope. The observed grain sizes are shown in Fig. 7. From the evaluation of the grain/particle statistics, we obtain average grain size of about 320 nm $\pm$ 60 nm and a particle size of 1.3 $\mu$m $\pm$ 0.23 $\mu$m.
To investigate cooperation of dynamic fields and AC-biasing, the sample is placed in a defocused 1030 nm beam while being kept in the focus of the 4500 nm pump beam. Images of the resulting PL were taken using a CMOS camera (Point Grey Research PGR-CM3-U3-50S5M-CS).
Figure 6: Abraded sample
Figure 7: a) Abraded sample as imaged by a SEM microscope with ‘particles’ and ‘grains’ highlighted in red and yellow, respectively; b) a schematic representation of a grain; c) grain size; d) particle size distributions.
Photoluminescence from CsPbBr3
To demonstrate tunnel ionization in another perovskite, we study the PL spectra of CsPbBr3 irradiated by a 4 µm laser, see Fig. 8. As for MAPbBr3, we observe that the PL intensity is exponentially sensitive to the strength of the external electric field. The crucial difference however is that our analytical predictions for MAPbBr3 do not describe the observed data in CsPbBr3. This could have been anticipated as in our derivations we relied on the cubic structure of the material. However, CsPbBr3 (unlike MAPbBr3) has an orthorhombic structure at room temperature so a more careful analysis must be in order. Further, we note that if the density of charge carriers is small (note that the intensity of the external electric field is relatively small in our experiment with CsPbBr3), then an electron can recombine only with its partner hole created in the tunnel ionization process or with localized trap states. These processes are not included in bimolecular radiative recombination used in our analysis of MAPbBr3.
To quantify the deviation from the $W_{fi}^{2}$-behavior observed in MAPbBr3 at small fields, we fit our CsPbBr3 data with $W_{fi}^{\delta}$, where $\delta$ is a fitting parameter. We find that $\delta\simeq 1.2$ can be used to describe the data. As this value is close to unity, we hypothesize that the radiative processes in CsPbBr3 for the considered parameters are mainly of monomolecular origin, see also Refs. Abiedh et al. (2021); Zhang et al. (2024).
[…]